Information about direct materials cost follows for a local company:

Standard price per materials gram $ 16
Actual quantity used 2,250 grams
Standard quantity allowed for production 2,375 grams
Price variance $ 14,400 F

Required:

What was the actual purchase price per gram? (Round your answer to 2 decimal places.)

Actual purchase price per gram
Answer #1

Price Variance = (SP-AP) * AQ

(16 - AP) * 2250 = 14,400

16 - AP = 6.4

AP = 9.60

Actual purchase price = 9.60 per gram